Based on a large sample of capacitors of a certain type, a 95% confidence interval for the mean capacitance, in μF, was computed to be (0.211, 0.265). Find a 90% confidence interval for the mean capacitance of this type of capacitor. Round the answers to three decimal places. The 90% confidence interval is ?
Answer #1

mean = (0.265 + 0.211)/2 = 0.238

ME(95) = (0.265 - 0.211)/2 = 0.027

ME(95) = 1.96*SE

ME(90) = 1.64*SE
ME(90) = 1.64*0.027/1.96 = 0.023

CI(90) = (0.238 - 0.023 , 0.238 + 0.023)
= (0.215 , 0.261)